# PRETRAINED CONFORMERS FOR AUDIO FINGERPRINTING AND RETRIEVAL

### v1

@@ Abstract
Conformers have shown great results in speech processing due to their ability to capture both local and global interactions. In this work, we utilize a self-supervised contrastive learning framework to train conformer-based encoders that are capable of generating unique embeddings for small segments of audio, generalizing well to previously unseen data. We achieve state-of-the-art results for audio retrieval tasks while using only 3 seconds of audio to generate embeddings. Our models are almost completely immune to temporal misalignments and achieve state-of-the-art results in cases of other audio distortions such as noise, reverb or extreme temporal stretching. Code and models are made publicly available and the results are easy to reproduce as we train and test using popular and freely available datasets of different sizes.
@@ 1 Introduction
Given a small excerpt of an audio, the task of content-based audio retrieval is to identify the original audio from which the excerpt was taken. Content-based audio retrieval is often performed using audio fingerprinting techniques; for every audio file of interest, a unique fingerprint is generated by taking smaller segments of the audio and generating low-dimensional representations of the segments. These low-dimensional representations collectively make up the audio fingerprint. Retrieval using an excerpt is performed by computing the low-dimensional representations of the excerpt and finding corresponding matches, often facilitated through vector search. Retrieval is limited to audio files for which fingerprints have been made and added to a database, while extending the database is performed by simply updating it with new fingerprints. Note that these low-dimensional representations do not necessarily need to be embeddings (or more generally vectors), instead, they can be as simple as pairs of time-frequency coordinates within the spectrogram, as is usually the case in more classical audio fingerprinting approaches.
The most common application of audio fingerprinting is in music identification services, such as those offered by Apple (Shazam) [1], Google (Now Playing [2]), SoundHound [3] and others. Audio fingerprinting is also used for detecting copyright infringements or tracking advertisements [1, 4]. In telecommunications, audio fingerprinting has been employed to search for duplicate audio when identifying unsolicited phone calls [5, 6] or even improve network performance by quickly identifying announcements played during session initiation [7, 8].
@@
Our contributions can be summarized as follows; we propose the first conformer-based architecture for neural audio fingerprinting, achieving state-of-the-art results for audio retrieval tasks. CNNs exploit local features effectively, while transformers are good at capturing content-based global interactions and have achieved state-of-the-art results in speech recognition [18, 19]. Conformers combine the best of both worlds, which makes them a suitable choice for audio fingerprinting tasks; when recognizing audio, the spectral content itself is just as important as its positioning in time within the audio. After an initial pretraining phase, our models learn to generalize well and can generate unique embeddings for previously unseen data. This paper is accompanied by three open source models and code which enables easy replication of our results. We highlight the importance of data augmentation during training, especially the task of generating difficult contrastive examples to prevent the models from overfitting.
@@ 2.2 Contrastive learning
Similar to previous work [2, 4, 9, 16], we use the SimCLR self-supervised contrastive learning framework [21] in which we wish to minimize the normalized temperature-scaled cross entropy loss (NT-Xent). The training process is visualized in Figure 2. Given a mini-batch of $B$ audio segments (samples), we generate a replica for each of the $B$ samples by applying various audio augmentation techniques. The original $z_{i}$} and replica $z_{j}$ are considered positive pairs, while the remaining 2(B1)2(B-1) samples are all treated as negative examples. In literature, $\mathrm{sim}(z_{i},z_{j})=z_{i}^{\top}z_{j}/\|z_{i}\|\|z_{j}\|$ is commonly used to denote the cosine similarity between two samples $z_{i}$ and $z_{j}$, which is equal to the dot product in our case, as we $\ell_{2}$ normalize the embeddings at the output of our model. The loss for a positive pair of samples $(i,j)$ is computed as:
@@
where $\mathbbm{1}_{[k\neq i]}\in\{0,1\}$ evaluates to $1$ iff $k\neq i$ and $\tau$ is the temperature parameter. The loss is now computed for each positive pair, both $(i,j)$ and $(j,i)$, and averaged into the total cost function. The training process is illustrated in Figure 2.
@@ 2.3 Data
Our models are trained using the Free Music Archive (FMA) [22]. FMA provides predefined subsets, categorized as small (8k tracks), medium (25k tracks), and large (106k tracks), with each subset already split into train-test-validation splits. Perhaps unconventionally, but we decided to train the small (1.5M), medium (8.8M) and large (26.2M) models using the small (8k), medium (25k) and large (106k) subsets of FMA, respectively. This choice was deliberate, as we wanted to make the small/medium models easier to reproduce. Although downsizing the models with fewer layers and parameters achieves this to some degree, a large dataset can still pose challenges for scientists lacking high-performance computing resources.
@@ 2.4 Generating positive pairs
We rely on the Python audiomentations library [23] to augment data; adding background noise at various SNR rates (datasets ESC-50 [24] and MS-SNSD [25]), adding colored noise at various SNR and decay rates, applying impulse response/reverb (dataset MIT IR Survey [26]), pitch-shifting the audio, applying polarity inversion, applying $\tanh$ distortion, time-stretching the audio and time-shifting the audio.
Despite using various audio augmentations, we report that SpecAugment [27] was crucial to prevent overfitting, especially for our large model.
@@ 2.4.1 Beta-distributed temporal shifting
In practice, it is desirable to allow for slight temporal mismatches when performing audio retrieval, i.e. the query segment does not have to be temporally aligned with the retrieved segment. Initially, we opted for a maximum of 5% mismatch ($150\text{\,}\mathrm{ms}$ or $2400\text{\,}\mathrm{s}$amples), thus we would randomly shift the replica (augmented audio) during training between 0 and 2400 samples.
Although no overfitting occurred, i.e. no substantial deviations between training and test/validation recall, the initial retrieval rates of models trained under such circumstances for relatively large time shifts, e.g. $120\text{\,}\mathrm{ms}$, were not satisfying.
After some investigation, similar to FaceNet [28], we concluded that it is important to train the model on what the authors refer to as hard examples. Thus, the training process was adjusted to encourage larger temporal shifts (i.e. harder examples) by sampling the time offset by which we shift audio from a beta distribution instead of a uniform distribution, as illustrated in Figure 3. Sampling from a beta distribution with parameters $\alpha=8$ and $\beta=2$ with the minimum and maximum offset values of $0\text{\,}\mathrm{ms}$ and $150\text{\,}\mathrm{ms}$ will on average apply an offset of $120\text{\,}\mathrm{ms}$. We report that this small change was crucial to improving performance and making our models almost completely robust to any temporal misalignments ranging from $0\text{\,}\mathrm{ms}$ to $150\text{\,}\mathrm{ms}$.
@@ 3 Experiments
Each result that is reported represents an average over 5 identical test runs. In a single test, we perform 5 queries for every audio file by selecting 5 linearly spaced 3-second long excerpts of each audio. As is commonly done in audio fingerprinting literature, we report on the top-1 and top-5 hit rate, often referred to as the exact and near matches or accuracy. For easier distinction, the top-5 hit rate will further be enlisted inside parentheses.
Fig. 3: Probability density function of the beta distribution for values $\beta=2$ and $\alpha=[5,6,7,8]$. Given $\alpha=8$ and a maximum value of 150, the mean would amount to 120.
@@ 3.1 Performance of conformer models
Of all distortions, our models are most robust to temporal shifts, which is supported by maintaining near-perfect results in the presence of such distortions, especially in terms of the top-5 hit rate. In Table 2, we provide the hit rates achieved when applying temporal shifts, expressed as a percentage of half the distance from the next fingerprint (i.e. $300\text{\,}\mathrm{ms}$). The models perform well on this type of distortion as they have learned to continuously decrease the similarity metric between two audio segments as the time distance between them increases, mimicking a monotonic function. Figure 4 illustrates this desirable property of the embeddings.
Table 2 provides the hit rates when dealing with reverb. Although these results are good, especially for the large conformer, the models are not as robust to reverb as to temporal offsets, which is further supported by Table 4 in which the overall performance for background and colored noises at various SNR levels is provided. Note that the results were evaluated using ESC-50 and the test split of MS-SNSD [24, 25], which none of the models has seen during training.
@@ 3.2.2 Noise distortions
In Table 3, we compare the performance of our best model with the current state-of-the-art approaches in noisy conditions. It is difficult to discern and point out a clear winner, since there are many variables at play, e.g. the used length of the query sample, the size of the embeddings or the subset of FMA [22]. The audio distortions are also not uniform, in NAFP [4] the results were reported under a varying SNR range, whereas in [15, 16, 17] independent test runs were repeated by fixing these values. The most important differences are listed under column Comments inside Table 3.
@@ 3.3 Discussion
For noisy audio, GraFPrint [16] claims to achieve the best results with a given query length of only 3 seconds, evaluated on the medium subset of FMA. NAFP is the most tested and cited implementation and has been confirmed to achieve similar results by several sources [4, 9, 16, 17]. Aside being evaluated on the large subset of FMA, another advantage of NAFP is that it achieves state-of-the-art performance in cases of extreme time stretching, while GraFPrint does not report on the results for such distortions.
Given the results from Table 3 and Figure 5, we conclude that our large conformer model (26.2M) matches the current state-of-the-art in terms of extreme temporal distortions (NAFP [4], PeakNetFP [9]) using a small query length of 3 seconds. Regarding noisy distortions, the large model has comparable performance to the state-of-the-art (GraFPrint [16]), especially at SNR rates of $10\text{\,}\mathrm{dB}$ and $15\text{\,}\mathrm{dB}$, while being the best performing model evaluated on the large subset of FMA [22]. Thus, our model achieves good results on both frontiers.

### v2

@@ Abstract
Conformers have shown great results in speech processing due to their ability to capture both local and global interactions. In this work, we utilize a self-supervised contrastive learning framework to train conformer-based encoders that are capable of generating unique embeddings for small segments of audio, which generalize well to previously unseen data. We obtain results comparable to existing approaches for audio retrieval tasks, while using only 3 seconds of audio to generate embeddings. Our models are almost completely immune to temporal misalignments and retain good performance in cases of audio distortions such as noise, reverb or extreme temporal stretching. Code and models are made publicly available and the results are easy to reproduce as we train and test using popular and freely available datasets of different sizes.
@@ 1 Introduction
Given a small excerpt of an audio file, the task of content-based audio retrieval is to identify the original file from which the excerpt was taken. Content-based audio retrieval is often performed using audio fingerprinting techniques. That is, for every audio file of interest, a unique fingerprint is generated by taking smaller segments of the file and generating low-dimensional representations of the segments. These low-dimensional representations collectively make up the audio fingerprint. Retrieval using an excerpt is performed by computing the low-dimensional representations of the excerpt and finding corresponding matches, often facilitated through vector search. Retrieval is limited to audio files for which fingerprints have been made and added to a database, while extending the database is performed by simply updating it with new fingerprints. Note that these low-dimensional representations do not necessarily need to be embeddings (or more generally vectors), instead, they can be as simple as time-frequency pairs within the spectrogram, as is usually the case in more classical audio fingerprinting approaches [1].
The most common application of audio fingerprinting is in music identification services, such as those offered by Apple (Shazam) [1], Google (Now Playing [2]), SoundHound [3] and others. Audio fingerprinting is also used for detecting copyright infringements or tracking advertisements [1, 4]. In telecommunications, audio fingerprinting has been employed to search for duplicate audio when identifying unsolicited phone calls [5, 6] or even to improve network performance by quickly identifying announcements played during session initiation [7, 8].
@@
Our contributions can be summarized as follows; we propose the first conformer-based architecture for neural audio fingerprinting, achieving competitive results for audio retrieval tasks. CNNs exploit local features effectively, while transformers are good at capturing content-based global interactions and have achieved state-of-the-art results in speech recognition [18, 19]. Conformers combine the best of both worlds, which makes them a suitable choice for audio fingerprinting tasks; when recognizing audio, the spectral content itself is just as important as its positioning in time within the audio. After an initial pretraining phase, our models learn to generalize well and can generate unique embeddings for previously unseen data. We highlight the importance of data augmentation during training, especially the task of generating difficult contrastive examples to prevent the models from overfitting. The paper is accompanied by three models of different sizes which we open source together with the code which enables easy replication of our results.
@@ 2.2 Contrastive learning
Similar to previous work [2, 4, 9, 16], we use the SimCLR self-supervised contrastive learning framework [21] in which we wish to minimize the normalized temperature-scaled cross entropy loss (NT-Xent). The training process is illustrated in Figure 2. Given a mini-batch of $B$ audio segments (samples), we generate a replica for each of the $B$ samples by applying various audio augmentation techniques. The original $z_{i}$} and replica $z_{j}$ are considered positive pairs, while the remaining 2(B1)2(B-1) samples are all treated as negative examples. In literature, $\mathrm{sim}(z_{i},z_{j})=z_{i}^{\top}z_{j}/\|z_{i}\|\|z_{j}\|$ is commonly used to denote the cosine similarity between two samples $z_{i}$ and $z_{j}$, which is equal to the dot product in our case, as we $\ell_{2}$ normalize the embeddings at the output of our model. The loss for a positive pair of samples $(i,j)$ is computed as:
@@
where $\mathbbm{1}_{[k\neq i]}\in\{0,1\}$ evaluates to $1$ iff $k\neq i$ and $\tau$ is the temperature parameter. The loss is now computed for each positive pair, both $(i,j)$ and $(j,i)$, and averaged into the total cost function.
@@ 2.3 Data
Our models are trained using the Free Music Archive (FMA) [22] which provides predefined data subsets of different sizes. We use the small subset (8k tracks) to train the small model (1.5M), medium (25k tracks) to train the medium model (8.8M), and large (106k tracks) for the large model (26.2M). Each of the subsets has an already predefined train-test-validation split that we follow.
@@ 2.4 Generating positive pairs
We rely on the Python audiomentations library [23] to augment data; adding background noise at various SNR rates (datasets ESC-50 [24] and MS-SNSD [25]), adding colored noise at various SNR and decay rates, applying impulse response/reverb (dataset MIT IR Survey [26]), pitch-shifting the audio, applying polarity inversion, applying $\tanh$ distortion, time-stretching the audio and time-shifting the audio. Despite using various audio augmentations, we observed that SpecAugment [27] was crucial to prevent overfitting, especially for our large model.
@@ 2.4.1 Beta-distributed temporal shifting
In practice, it is desirable to allow for slight temporal mismatches when performing audio retrieval, i.e., the query segment does not have to be temporally aligned with the retrieved segment. Initially, we opted for a maximum of 5% mismatch ($150\text{\,}\mathrm{ms}$ or $2400\text{\,}\mathrm{s}$amples), thus we would randomly shift the replica (augmented audio) during training between 0 and 2400 samples. Although no overfitting occurred, that is, no substantial deviations between training and test/validation recall, the initial retrieval rates of models trained under such circumstances for relatively large time shifts, e.g. $120\text{\,}\mathrm{ms}$, were not satisfying. Simply put, our models were underfit for tasks with large time shifts.
After some investigation, similar to FaceNet [28], we concluded that it is important to train the model on what the authors refer to as hard examples. Thus, the training process was adjusted to encourage larger temporal shifts (harder examples) by sampling the time offset by which we shift audio from a beta distribution instead of a uniform distribution. As illustrated in Figure 3, sampling from a beta distribution with parameters $\alpha=8$ and $\beta=2$ with the minimum and maximum offset values of $0\text{\,}\mathrm{ms}$ and $150\text{\,}\mathrm{ms}$ will on average apply an offset of $120\text{\,}\mathrm{ms}$. This seemingly small change was crucial to improving performance and making our models almost completely robust to any temporal misalignments ranging from $0\text{\,}\mathrm{ms}$ to $150\text{\,}\mathrm{ms}$.
@@ 3 Experiments
All reported results represent an average over 5 identical test runs. In a single test, we perform 5 queries for every audio file by selecting 5 linearly spaced 3-second long excerpts of each audio. As is commonly done in audio fingerprinting literature, we report on the top-1 and top-5 hit rate, often referred to as the exact and near matches or accuracy [4, 16, 17]. For easier distinction, the top-5 hit rate will further be enlisted inside parentheses.
Fig. 3: Probability density function of the beta distribution for values $\beta=2$ and $\alpha=[5,6,7,8]$. Given $\alpha=8$ and a maximum value of $150\text{\,}\mathrm{ms}$, the mean would amount to $120\text{\,}\mathrm{ms}$.
@@ 3.1 Performance of conformer models
Of all distortions, our models are most robust to temporal shifts, which is supported by maintaining near-perfect results in the presence of such distortions, especially in terms of the top-5 hit rate. In Table 2, we provide the hit rates achieved when applying temporal shifts, expressed as a percentage of half the distance from the next fingerprint (i.e. $300\text{\,}\mathrm{ms}$). Our models perform well in this type of distortion as they have learned to continuously decrease the similarity metric between two audio segments as the time distance between them increases, mimicking a monotonic function. Figure 4 illustrates this desirable property of the embedding’s robustness to temporal misalignments.
In Table 2 we further provide the hit rates when dealing with reverb. Although these results are good, especially for the large conformer, the models are not as robust to reverb as to temporal offsets, which is further supported by Table 4 in which the overall performance for background and colored noises at various SNR levels is provided. Note that the results were evaluated using ESC-50 and the test split of MS-SNSD [24, 25], which none of the models had seen during training.
@@ 3.2.2 Noise distortions
In Table 3, we compare the performance of our best model with the current state-of-the-art approaches in noisy conditions. It is difficult to discern and point out a clear winner, since there are many variables at play, e.g., the length of the query sample used, the size of the embeddings or the subset of FMA [22]. The audio distortions are also not uniform, in NAFP [4] the results were reported under a varying SNR range, whereas in [15, 16, 17] independent test runs were repeated by fixing these values. The most important differences are listed under column Comments inside Table 3.
@@ 3.3 Discussion
For noisy audio, GraFPrint [16] claims to achieve the best results with a given query length of only 3 seconds, evaluated on the medium subset of FMA. NAFP is the most referenced and tested implementation and has been confirmed to achieve similar results by several sources [4, 9, 16, 17]. Aside being evaluated on the large subset of FMA, another advantage of NAFP is that it achieves state-of-the-art performance in cases of extreme time stretching, while GraFPrint does not report on the results for such distortions.
Given the results from Table 3 and Figure 5, we conclude that our large conformer model (26.2M) is comparable to the current state-of-the-art in terms of extreme temporal distortions (NAFP [4], PeakNetFP [9]) using a small query length of 3 seconds. Regarding noisy distortions, the large model has comparable performance to GraFPrint [16], especially at SNR rates of $10\text{\,}\mathrm{dB}$ and $15\text{\,}\mathrm{dB}$, while being the best performing model evaluated on the large subset of FMA [22]. Thus, our model achieves good results on both frontiers.